# Dual-Head Knowledge Distillation: Enhancing Logits Utilization with an Auxiliary Head

### v1

@@ B.3 Propositions for the Modified Loss
$\mathcal{L}_{\mathrm{overall}}=\mathcal{L}_{CE}+\alpha\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}},$ (15)
@@
$\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}=\tau^{2}\sum\limits_{i=1}^{B}\sum\limits_{k=1}^{K}\mathcal{KL}\left(\left.\left[\frac{1}{2},\frac{1}{2}\right]\right\|\left[\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right),1-\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right)\right]\right),$
as defined in Equation 5.
@@
$\left[(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})\right]^{\top}\bm{w}_{k}\leq 0.$ (17)
@@
$\left[(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})\right]^{\top}\bm{w}_{k}\\ =(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))\left[(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})^{\top}\bm{w}_{k}\right]\leq 0.$
@@
$\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{w}_{k}}=-(\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}})-\alpha\bm{F}_{\bm{w}\mathrm{-obstacle}}^{\mathrm{BinaryKL\text{-}Norm}},$ (18)
@@
$\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}=\sum\limits_{i=1}^{n_{k}}(1-p_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$ $\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k^{\prime}\neq k}^{K}\sum\limits_{j=1}^{n_{k^{\prime}}}p_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},i})\bm{h}^{\mathcal{S}}_{k^{\prime},j},$ (19)
@@
$-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}}{\partial\bm{w}_{k}}=\sum_{k^{\prime}=1}^{K}\sum_{j=1}^{n_{k^{\prime}}}\tau(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})$ (20)
@@
Based upon Equation 9, 20, we can prove Proposition 3.
@@
$\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{h}}=-(\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}})-(\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}),$ (21)
@@
$\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}=(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},$ $\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}}=\tau(\frac{1}{2}-w_{c}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{c},$ (22)
$\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k},$ $\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}=\tau\sum\limits_{k\neq c}^{K}(w_{k}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}})-\frac{1}{2})\bm{w}^{\mathcal{S}}_{k}.$
@@
$-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}}{\partial\bm{w}_{k}}=\underbrace{\tau(\frac{1}{2}-w_{c}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}}}+\underbrace{\left(-\tau\sum\limits_{k\neq c}^{K}(w_{k}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}})-\frac{1}{2})\bm{w}^{\mathcal{S}}_{k}\right)}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}}.$ (23)
Based upon Equation 13, 23, we can prove Proposition 4.
@@ Appendix D Comparison between BinaryKL and BinaryKL-Norm
Table 8: Comparison between BinaryKL and BinaryKL-Norm. For the BinaryKL loss, we compare the results with a fixed $\tau=2$ and the best $\tau$. Teachers and students are in the same architectures. All the experiments are done on the CIFAR-100 dataset.
Teacher resnet56 resnet110 resnet32$\times$4 WRN-40-2 WRN-40-2 VGG13
Student resnet20 resnet32 resnet8$\times$4 WRN-16-2 WRN-40-1 VGG8
DHKD using BinaryKL with $\tau=2$ 68.95 72.64 76.56 75.24 73.24 74.26
DHKD using BinaryKL with the best $\tau$ 70.23 73.11 76.86 76.18 74.15 74.78
DHKD using BinaryKL-Norm with $\tau=2$ 71.19 73.92 76.54 76.36 75.25 74.84
Table 9: Comparison between BinaryKL and BinaryKL-Norm. For the BinaryKL loss, we compare the results with a fixed $\tau=2$ and the best $\tau$. Teachers and students are in different architectures. All the experiments are done on the CIFAR-100 dataset.
Teacher resnet32$\times$4 WRN-40-2 VGG13 ResNet-50 resnet32$\times$4
Student ShuffleNet-V1 ShuffleNet-V1 MBN-V2 MBN-V2 ShuffleNet-V2
DHKD using BinaryKL with $\tau=2$ 75.60 75.92 69.85 69.54 77.63
DHKD using BinaryKL with the best $\tau$ 77.04 76.67 70.03 70.66 77.82
DHKD using BinaryKL-Norm with $\tau=2$ 76.78 77.25 70.09 71.08 77.99
@@ Appendix F Implementation Details
We perform experiments on two benchmark datasets: CIFAR-100 (Krizhevsky & Hinton, 2009) and ImageNet (Deng et al., 2009). CIFAR-100 covers $100$ categories. It contains $50000$ images in the train set and $10000$ images in the test set. ImageNet covers $1000$ categories of images. It contains $1.28$ million images in the train set and $50000$ images in the test set.
@@
Appendix G t-SNE Visualization of the Features
(a) Vanilla
(b) KD
(c) DHKD
(d) Teacher
Figure 5: t-SNE visualization of features learned by different methods. We do the visualization on the test set of CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
t-SNE (t-distributed Stochastic Neighbor Embedding) (Van der Maaten & Hinton, 2008) is a famous unsupervised non-linear dimensionality reduction technique for visualizing high-dimensional data. We use t-SNE to visualize the features (the outputs of the models’ backbones) learned by different models: vanilla student without distillation, models trained by KD, DHKD, and the teacher model.
Appendix H Difference of Correlation Matrices
(a) Vanilla
(b) DKD
(c) DHKD Aux
(d) DHKD
Figure 6: The difference between the correlation matrices of the teacher’s and student’s logits. “DHKD Aux” is short for DHKD’s auxiliary head. We do the visualization on the test set of CIFAR-100, with resnet32$\times$4 as the teacher and resnet8$\times$4 as the student. The performance sorting of four classifier heads is DHKD (76.54) $>$ DKD (76.37) $>$ DHKD’s Aux (76.20) $>$ Vanilla (72.17).
Appendix I Training Time
@@
We also recorded the training time consumption on ImageNet. As stated in Section F, we utilize the gradient alignment module only to achieve state-of-the-art performance on CIFAR-100 and do not apply it to ImageNet. Consequently, our method only needs 616 ms for each batch, nearly equal to DKD which costs 614 ms.
Table 10: Training time (per batch) and the numbers of extra parameters on CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
Method KD RKD FitNet OFD CRD ReviewKD DKD DHKD w/o GA DHKD w/ GA
Top-1 Accuracy 73.33 71.90 73.50 74.95 75.51 75.63 76.32 76.16 76.54
Time (ms) 11 25 14 19 41 26 11 14 21
# params 0 0 16.8k 86.9k 12.3M 1.8M 0 25.6k 25.6k
Appendix J Parameter Sensitivity Analysis
@@
Table 11: The parameter sensitivity analysis on CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
$\alpha$ 0.2 0.5 1 2 5
Top-1 Accuracy 75.66 75.85 76.24 76.54 67.56
Appendix K The Performance of “dual head + vanilla KD”
@@
Table 12: Comparison between DHKD and “dual head + vanilla KD”.
Teacher Student without distillation vanilla KD dual head + vanilla KD DHKD
WRN-40-2 WRN-16-2 73.26 74.92 75.24 76.36
ResNet50 MBN-V2 64.60 67.35 68.76 71.08

### v2

@@ B.3. Propositions for the Modified Loss
(19) $\mathcal{L}_{\mathrm{overall}}=\mathcal{L}_{CE}+\alpha\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}},$
@@
$\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}=\tau^{2}$ $\sum\limits_{i=1}^{B}\sum\limits_{k=1}^{K}\mathcal{KL}\Big{(}\left.\left[\frac{1}{2},\frac{1}{2}\right]\right\|$
(20) $\left[\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right),1-\sigma\left(\frac{z_{i,k}^{\mathcal{S}}-z_{i,k}^{\mathcal{T}}}{\tau}\right)\right]\Big{)},$
as defined in Equation 3.3.
@@
(22) $\left[(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})\right]^{\top}\bm{w}_{k}\leq 0.$
@@
$\left[\left(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j})\right)\left(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j}\right)\right]^{\top}$ $\bm{w}_{k}$
(23) $=(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))\Big{[}(\bm{h}^{\mathcal{S}}_{k^{\prime},j}$ $-\bm{h}^{\mathcal{T}}_{k^{\prime},j})^{\top}\bm{w}_{k}\Big{]}\leq 0.$
@@
(24) $\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{w}_{k}}=-(\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}})-\alpha\bm{F}_{\bm{w}\mathrm{-obstacle}}^{\mathrm{BinaryKL\text{-}Norm}},$
@@
(25) $\bm{F}_{\bm{w}\mathrm{\text{-}pull}}^{\mathrm{CE}}=\sum\limits_{i=1}^{n_{k}}(1-p_{k}(\bm{h}^{\mathcal{S}}_{k,i}))\bm{h}^{\mathcal{S}}_{k,i},$ $\bm{F}_{\bm{w}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k^{\prime}\neq k}^{K}\sum\limits_{j=1}^{n_{k^{\prime}}}p_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},i})\bm{h}^{\mathcal{S}}_{k^{\prime},j},$
@@
(26) $-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}}{\partial\bm{w}_{k}}=\sum_{k^{\prime}=1}^{K}\sum_{j=1}^{n_{k^{\prime}}}\tau(\frac{1}{2}-w_{k}(\bm{h}^{\mathcal{S}}_{k^{\prime},j},\bm{h}^{\mathcal{T}}_{k^{\prime},j}))(\bm{h}^{\mathcal{S}}_{k^{\prime},j}-\bm{h}^{\mathcal{T}}_{k^{\prime},j})$
@@
Based upon Equation 11, 26, we can prove Proposition 4.
@@
(27) $\frac{\partial\mathcal{L}_{\mathrm{overall}}}{\partial\bm{h}}=-(\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}})-(\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}+\alpha\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}),$
@@
(28) $\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{CE}}=(1-p_{c}(\bm{h}^{\mathcal{S}}))\bm{w}^{\mathcal{S}}_{c},$
$\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}}=\tau(\frac{1}{2}-w_{c}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{c},$
$\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{CE}}=-\sum\limits_{k\neq c}^{K}p_{k}(\bm{h}^{\mathcal{S}})\bm{w}^{\mathcal{S}}_{k},$
$\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}=\tau\sum\limits_{k\neq c}^{K}(w_{k}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}})-\frac{1}{2})\bm{w}^{\mathcal{S}}_{k}.$
@@
$-\frac{\partial\mathcal{L}_{\mathrm{BinaryKL\text{-}Norm}}}{\partial\bm{w}_{k}}=$ $\underbrace{\tau(\frac{1}{2}-w_{c}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}}))\bm{w}^{\mathcal{S}}_{c}}_{\bm{F}_{\bm{h}\mathrm{\text{-}pull}}^{\mathrm{BinaryKL\text{-}Norm}}}+$
(29) $\underbrace{\left(-\tau\sum\limits_{k\neq c}^{K}(w_{k}(\bm{h}^{\mathcal{S}},\bm{h}^{\mathcal{T}})-\frac{1}{2})\bm{w}^{\mathcal{S}}_{k}\right)}_{\bm{F}_{\bm{h}\mathrm{\text{-}push}}^{\mathrm{BinaryKL\text{-}Norm}}}.$
Based upon Equation 16, B.3, we can prove Proposition 5.
@@ Appendix D Comparison between BinaryKL and BinaryKL-Norm
Table 8. Comparison between BinaryKL and BinaryKL-Norm. For the BinaryKL loss, we compare the results with a fixed $\tau=2$ and the best $\tau$. Teachers and students are in the same architectures. All the experiments are done on the CIFAR-100 dataset.
[…]
. Teacher resnet56 resnet110 resnet32$\times$4 WRN-40-2 WRN-40-2 VGG13 Student resnet20 resnet32 resnet8$\times$4 WRN-16-2 WRN-40-1 VGG8 DHKD using BinaryKL with $\tau=2$ 68.95 72.64 76.56 75.24 73.24 74.26 DHKD using BinaryKL with the best $\tau$ 70.23 73.11 76.86 76.18 74.15 74.78 DHKD using BinaryKL-Norm with $\tau=2$ 71.19 73.92 76.54 76.36 75.25 74.84
Table 9. Comparison between BinaryKL and BinaryKL-Norm. For the BinaryKL loss, we compare the results with a fixed $\tau=2$ and the best $\tau$. Teachers and students are in different architectures. All the experiments are done on the CIFAR-100 dataset.
[…]
. Teacher resnet32$\times$4 WRN-40-2 VGG13 ResNet-50 resnet32$\times$4 Student ShuffleNet-V1 ShuffleNet-V1 MBN-V2 MBN-V2 ShuffleNet-V2 DHKD using BinaryKL with $\tau=2$ 75.60 75.92 69.85 69.54 77.63 DHKD using BinaryKL with the best $\tau$ 77.04 76.67 70.03 70.66 77.82 DHKD using BinaryKL-Norm with $\tau=2$ 76.78 77.25 70.09 71.08 77.99
Table 10. Training time (per batch) and the numbers of extra parameters on CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
[…]
. Method KD RKD FitNet OFD CRD ReviewKD DKD DHKD w/o GA DHKD w/ GA Top-1 Accuracy 73.33 71.90 73.50 74.95 75.51 75.63 76.32 76.16 76.54 Time (ms) 11 25 14 19 41 26 11 14 21 # params 0 0 16.8k 86.9k 12.3M 1.8M 0 25.6k 25.6k
Table 11. The parameter sensitivity analysis on CIFAR-100. We set resnet32$\times$4 as the teacher and resnet8$\times$4 as the student.
[…]
. $\alpha$ 0.2 0.5 1 2 5 Top-1 Accuracy 75.66 75.85 76.24 76.54 67.56
Table 12. Comparison between DHKD and “dual head + vanilla KD”.
[…]
. Teacher Student without distillation vanilla KD dual head + vanilla KD DHKD WRN-40-2 WRN-16-2 73.26 74.92 75.24 76.36 ResNet50 MBN-V2 64.60 67.35 68.76 71.08
Table 13. Comparison of mIoU performance on Cityscapes with different distillation methods.
Method Teacher Student SKD IFVD CWD CIRKD DIST DHKD
mIoU (%) 78.07 74.21 75.42 75.59 75.55 76.38 77.10 77.13
@@
  • SimKD (Chen et al., 2022) reuses the teacher’s classifier and aligns student features with a single L2 loss, achieving teacher-level performance with broad architectural compatibility;
  • CAT-KD (Guo et al., 2023) enhances both interpretability and performance by distilling class activation maps, improving the student’s ability to identify class-discriminative regions for classification;
@@ Appendix F Implementation Details
We perform experiments on two benchmark datasets: CIFAR-100 (Krizhevsky and Hinton, 2009) and ImageNet (Deng et al., 2009). CIFAR-100 covers $100$ categories. It contains $50000$ images in the train set and $10000$ images in the test set. ImageNet covers $1000$ categories of images. It contains $1.28$ million images in the train set and $50000$ images in the test set.
@@
Appendix G Training Time
@@
As stated in Section F, we utilize the gradient alignment module only to achieve state-of-the-art performance on CIFAR-100 when using DHKD alone. We do not apply it when testing on ImageNet and testing with ReviewKD on CIFAR-100. As a result, in most cases, we would not use the gradient alignment module. We recorded the training time consumption on ImageNet. The results show that our method only needs 616 ms for each batch, nearly equal to DKD which costs 614 ms, demonstrating the high efficiency of our DHKD methed.
Appendix H Parameter Sensitivity Analysis
@@
Appendix I The Performance of “dual head + vanilla KD”
@@
Appendix J Experiments on Semantic Segmentation
We conduct additional experiments on semantic segmentation, evaluating DHKD on the Cityscapes dataset (cordts2016cityscapes) with DeepLabV3 (chen2017deeplab) (ResNet-101 teacher and ResNet-18 student). Distillation is applied only to the final class predictions. As summarized in Table 13, DHKD achieves competitive performance compared to existing methods, confirming its effectiveness. Additionally, DHKD’s modular design ensures adaptability to various backbones, supporting teacher-student pairs with different architectures (e.g., ResNet to MBN). Extending DHKD to more diverse configurations, such as CNN-to-Transformer setups, presents an exciting direction for future work.